# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ 5 Shaving-Off Bits with the Eigenvalue Method
$\beta_{n-1}\cdot x_{n-1}=\alpha_{n}^{2}+\tilde{\mathcal{A}}_{i}(x_{1},\dots,x_{n-2},\alpha_{n}).$ (59)
(If $\beta_{n-1}=0$, then we simply rename the variables.) We can then further eliminate the variable $x_{n-1}$ in the polynomial system $\mathcal{F}\setminus\{f_{n-1},f_{n}\}\subset\mathbb{F}_{q}[x_{1},\dots,x_{n}]$ by substituting the previous affine equation for $x_{n-1}$ and $x_{n}=\alpha_{n}$. I.e., by finding one eigenvalue we can eliminate two variables. We also note that the Hydra DRL Gröbner basis (Section 4.2) satisfies the shape of Equation 58. Then, by iterating this variable elimination Equation 56 improves to
$\mathcal{O}\left(\begin{rcases}\begin{dcases}\sum_{i=1}^{\frac{n}{2}}N^{n-2\cdot i}\cdot\Big{(}2^{\omega\cdot(2\cdot i-1)}+D\big{(}2^{2\cdot i},q\big{)}\Big{)},&2\mid n,\\ \sum_{i=1}^{\frac{n-1}{2}}N^{n-1-2\cdot i}\cdot\Big{(}2^{\omega\cdot 2\cdot i}+D\big{(}2^{2\cdot i+1},q\big{)}\Big{)},&2\nmid n\end{dcases}\end{rcases}\right)$ (60)
$=\mathcal{O}\left(\begin{rcases}\begin{dcases}2^{\omega}\cdot\frac{2^{\omega\cdot n}-N^{n}}{2^{2\cdot\omega}-N^{2}}+\sum_{i=1}^{\frac{n}{2}}N^{n-2\cdot i}\cdot D\big{(}2^{2\cdot i},q\big{)},&2\mid n,\\ 2^{2\cdot\omega}\cdot\frac{2^{\omega\cdot(n-1)}-N^{n}}{2^{2\cdot\omega}-N^{2}}+\sum_{i=1}^{\frac{n-1}{2}}N^{n-1-2\cdot i}\cdot D\big{(}2^{2\cdot i+1},q\big{)},&2\nmid n\end{dcases}\end{rcases}\right),$ (61)
where we again assume that every characteristic polynomial has at most $N$ roots in $\mathbb{F}_{q}$.
Again for $N=1$, this implies that
$\log_{2}\left(\begin{rcases}\begin{dcases}2^{\omega}\cdot\frac{2^{\omega\cdot n}-1}{2^{2\cdot\omega}-1},&2\mid n,\\ 2^{2\cdot\omega}\cdot\frac{2^{\omega\cdot(n-1)}-1}{2^{2\cdot\omega}-1},&2\nmid n\end{dcases}\end{rcases}\right)$ $\leq\omega\left(n+1\right)-\log_{2}\left(2^{2\cdot\omega}-1\right)$ (62)
$\approx\begin{dcases}2\cdot n-1.9069,&\omega=2,\\ 3\cdot n-2.9773,&\omega=3.\end{dcases}$ (63)
I.e., we expect an improvement of $\approx\omega$ compared to generic probabilistic term order conversion.
6 Boolean Macaulay Matrices
Let $\mathcal{F}\subset\mathbb{F}_{2}[x_{1},\dots,x_{n}]$ be a polynomial system. If we are only interested in $\mathbb{F}_{2}$-valued solutions, then it is sufficient to compute a Gröbner basis of the ideal $\left(\mathcal{F},x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$. On the other hand, via the field equations $F=\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ we can reduce any monomial in the Macaulay matrix to a square-free monomial. Effectively, this reduces the full Macaulay matrix to a square-free submatrix, the so-called Boolean Macaulay matrix. Gaussian elimination for Boolean DRL Macaulay matrices over $\mathbb{F}_{2}$ was analyzed in [BFSS13] by Bardet et al.
Interestingly, for Ciminion & Hydra we found zero-dimensional DRL Gröbner bases with leading monomials $x_{i}^{2}$. Hence, for Gröbner basis computations we can use these bases to reduce the Macaulay matrix to a “Boolean” one over an arbitrary finite field $\mathbb{F}_{q}$. This motivates the following definitions.
Definition 6.1.
  1. (i)
  2. (ii)
    $\mathcal{F}$ is a $>$-Gröbner basis.
Definition 6.2 (cf. [BFSS13, Definition 1]).
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. For the (inhomogeneous) $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$:
  • Columns are indexed by square-free monomials $s\in P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(m\right)\leq d$ sorted via $>$.
  • Rows are index by polynomials $t\cdot f$ where $t\in P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ is a square-free monomial and $f\in\mathcal{F}$ such that $\deg\left(t\cdot f\right)\leq d$.
Note that any monomial $t\in P$ can be reduced to a square-free monomial $\hat{t}\equiv t\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with $\deg\left(\hat{t}\right)\leq\deg\left(t\right)$ due to degree compatibility of $>$. Since remainders with respect to a Gröbner basis are unique, see [CLO15, Chapter 2 §6 Proposition 1], for the remainder of $t\cdot f$ it does not matter whether we reduce $t$ first and then $\hat{t}\cdot f$ or $t\cdot f$ at once. So it is indeed sufficient to consider only products with square-free monomials.
Obviously, the remainder $r_{t\cdot f}=t\cdot f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ is polynomial consisting of square-free terms. Moreover, by elementary properties of multivariate division by remainder with respect to a degree compatible term order, see [CLO15, Chapter 2 §3 Theorem 3], we have that
$t\cdot f=\sum_{i=1}^{m}g_{i}\cdot f_{i}^{\text{Bool}}+r_{t\cdot f},$ (64)
where $\deg\left(g_{i}\cdot f_{i}^{\text{Bool}}\right),\deg\left(r_{t\cdot f}\right)\leq\deg\left(t\cdot f\right)$. I.e., the $K$-Boolean Macaulay matrix can be constructed via row operations from the Macaulay matrix $M_{\leq d}$ of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$. Alternatively, we can obtain the $K$-Boolean Macaulay matrix by pivoting non-square-free monomials in $M_{\leq d}$ via rows coming from $\mathcal{F}_{\text{Bool}}$, and then performing an auxiliary Gaussian elimination step that stops after all redundant pivots coming from $\mathcal{F}_{\text{Bool}}$ have been eliminated.
Let $M_{\leq d}$ be a Macaulay matrix, we denote with $\operatorname{rowsp}\left(M_{\leq d}\right)$ the row space basis, i.e. the polynomial system obtained via Gaussian elimination on $M_{\leq d}$. Analog, we use this notation for Boolean Macaulay matrices.
Definition 6.3 (Boolean Solving Degree).
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,\allowbreak x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. The Boolean solving degree of $\mathcal{F}\cup\mathcal{F}_{\text{Bool}}$ is the least degree $d$ such that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ contains a Gröbner basis with respect to $>$. We denote it by $\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)$.
Note that $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)$ might not contain zero-dimensional Gröbner basis of the full system. A trivial example is =$\mathcal{F}=\left\{x_{1}^{2},x_{1}\cdot x_{2},x_{2}^{2}\right\}$. Therefore, to extract the Gröbner basis from $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ we set up the ideal of $>$-leading terms and then compute its minimal generating set. With the minimal generating set we can then filter the Gröbner basis. (It is also obvious that all non-square-free monomials are generated by the leading terms of $\mathcal{F}_{\text{Bool}}$, hence it is indeed sufficient to just consider $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$.)
Theorem 6.4.
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}},\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be polynomial systems, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. Then
  1. (1)
  2. (2)
    If $d=\operatorname{sd}^{\text{Bool}}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)<n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires
    $\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\left(\sum_{i=1}^{d}\binom{n}{i}\right)^{\omega}\right)$
    field operations.
  3. (3)
    If $d=\operatorname{sd}_{>}\left(\mathcal{F}\cup\mathcal{F}_{\text{Bool}}\right)\geq n$, then Gaussian elimination on $M_{\leq d}^{\text{Bool}}$ requires at most
    field operations to produce a $>$-Gröbner basis.
Proof.
For (1), we pivot non-square-free monomials via polynomials from $\mathcal{F}_{\text{Bool}}$, then we perform the auxiliary Gaussian elimination step that produces $M_{\leq d}^{\text{Bool}}$ as submatrix, and finally we perform Gaussian elimination on $M_{\leq d}^{\text{Bool}}$. Since Gaussian elimination does not depend on the order of reductions $\operatorname{rowsp}\left(M_{\leq d}^{\text{Bool}}\right)\cup\mathcal{F}_{\text{Bool}}$ has to contain the $>$-Gröbner basis.
For (2), trivially in degree $i$ there exist $\binom{n}{i}$ square-free monomials, and every $f\in\mathcal{F}$ can contribute to at most $\binom{n}{i}$ rows. I.e., the number of columns is bounded by $D=\sum_{i=1}^{d}\binom{n}{i}$, and the number of rows by $\left|\mathcal{F}\right|\cdot D$. The complexity of Gaussian elimination then follows from [Sto00, §2.2].
For (3), let $s\in P$ be a non-square-free monomial with $\deg\left(s\right)>n$, then $\hat{s}\equiv s\mod\left(\mathcal{F}_{\text{Bool}}\right)$ where $\hat{s}\in P$ is square-free, in particular $\deg\left(\hat{s}\right)\leq n$. Therefore, is suffices to perform Gaussian elimination on the full $K$-Boolean Macaulay matrix. Moreover, it is well known that $\sum_{i=0}^{n}\binom{n}{i}=2^{n}$. ∎
Remark 6.5.
This theorem can be refined by replacing the binomial coefficient $\binom{n}{i}$ by the number of monomials in $P\setminus\operatorname{in}_{>}\left(\mathcal{F}_{\text{Bool}}\right)$ of degree $i$.
So far our analysis does not differ from the one over $\mathbb{F}_{2}$, however the binary field has a special feature. Division by remainder with respect to $\left(x_{1}^{2}-x_{1},\dots,x_{n}^{2}-x_{n}\right)$ is “for free”, i.e. its complexity is negligible compared to Gaussian elimination. For general $K$-Boolean polynomial systems this is not necessarily the case. Luckily, the complexity of general division by remainder with respect to quadratic $K$-Boolean polynomial systems follows as corollary to [Ten19, Proposition 3.10].
Corollary 6.6.
Let $K$ be a field, let $\mathcal{F}_{\text{Bool}}\subset P=K[x_{1},\dots,x_{n}]$ be a quadratic polynomial system, let $f\in P$ be a polynomial, and let $>$ be a degree compatible term order on $P$. Assume that $\mathcal{F}_{\text{Bool}}$ is $K$-Boolean with respect to $>$. Then computation of $r_{f}=f\mod\left(\mathcal{F}_{\text{Bool}}\right)$ with respect to $>$ requires
@@ 6 Boolean Macaulay Matrices
$\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}\right)$
@@
$\in\mathcal{O}\left(\frac{n^{2}+3\cdot n+2}{2}\cdot\deg\left(f\right)\cdot\binom{n+\deg\left(f\right)-1}{\deg\left(f\right)}\right)$
field operations.
Proof.
There exist $\sum_{i=0}^{\deg\left(f\right)}\binom{n+i-1}{i}$ monomials of degree $\leq\deg\left(f\right)$ and $\binom{n+1}{2}+\binom{n}{1}+1$ monomials of degree $\leq 2$. Now the claim follows from [Ten19, Proposition 3.10]. ∎
Thus, the complexity of constructing the $K$-Boolean Macaulay matrix $M_{\leq d}^{\text{Bool}}$ with respect to quadratic $K$-Boolean system $\mathcal{F}_{\text{Bool}}$ is bounded by
$\mathcal{O}\left(\left|\mathcal{F}\right|\cdot\frac{n^{2}+3\cdot n+2}{2}\cdot\left(\sum_{i=0}^{d}\binom{n+i-1}{i}\right)\cdot\left(\sum_{i=0}^{d}\binom{n}{i}\right)\right).$ (65)
Moreover, we have the inequality
[…]
on M_≤d^Bool
[…]
on M_≤d
[…]
on M_≤d^Bool
[…]
on M_≤d
[…]
I.e. for $\omega=2$, if we ignore implied constants and prefactors, then we expect that construction of the $K$-Boolean Macaulay matrix dominates the Gröbner basis computation for general quadratic $K$-Boolean polynomial systems.
Of course, the $K$-Boolean Macaulay matrix generalizes to more advanced linear algebra-based Gröbner basis algorithms. Current state-of-the-art are Faugère’s F4 [Fau99] and Matrix-F5 [Fau02] algorithms. These algorithms perform Gaussian elimination on the Macaulay matrix $M_{\leq d}$ for increasing values of $d$. Moreover, F4 and F5 avoid redundant rows via efficient selection criteria. We illustrate a simplified variant of F4. Let $\mathcal{F}\subset P=K[x_{1},\dots,x_{n}]$ be a polynomial system, let $>$ be a term order, and let $M_{\leq d}$ be the Macaulay matrix for $\mathcal{F}$ in degree $d$. Then, we perform Gaussian elimination on $M_{\leq d}$ which yields the row space basis $\mathcal{G}_{1}$, now we set up a new Macaulay matrix $M_{\leq d}$ for $\mathcal{G}_{1}$ and again compute the row space basis $\mathcal{G}_{2}$ via Gaussian elimination. We iterate this procedure until the row space basis becomes stationary, i.e. $\mathcal{G}_{i}=\mathcal{G}_{i+1}$, then we verify whether $\mathcal{G}_{i}$ is already a $>$-Gröbner basis. If not then we pass to degree $d+1$ and repeat the procedure. Obviously, if the input system contains a $K$-Boolean polynomial system, then we can replace every full Macaulay matrix by its $K$-Boolean submatrix.
7 Cryptanalysis of Ciminion & Hydra
In this section we discuss how our results from the previous sections can be applied to cryptanalysis of Ciminion and Hydra.
7.1 Ciminion
For Gröbner basis cryptanalysis of Ciminion [DGGK21, § 4.4] the designers considered a weaker scheme where the keys are added after the permutations $p_{E}$ instead of before the rolling functions, see [DGGK21, Fig. 7]. The designer claim that the weaker scheme is simpler to analyze than Ciminion. We reject this claim for the iterated Ciminion polynomial system, after all we have computed a DRL Gröbner basis for Ciminion with rather simple affine variable eliminations, see LABEL:Th:_Ciminion_Gr\"obner_basis.
The designers assumed that weakened Ciminion polynomial systems are regular, moreover for regular systems it is well-known that the degree of regularity is equal to the Macaulay bound. Then, the DRL Gröbner basis complexity is bounded by [BFS04, Theorem 7]
$\mathcal{O}\left(m\cdot d_{\operatorname{reg}}\cdot\binom{n+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right).$ (67)
For Ciminion cryptanalysis, it is then assumed that the DRL Gröbner basis computation is the most costly part in an attack. The Ciminion designers did not provide a dedicated analysis for the iterated polynomial model from Definition 2.1. They mostly analyze polynomial models where $p_{E}$ and $p_{C}$ or respectively $p_{E}\circ p_{C}$ are modeled without intermediate variables, i.e. they used a single polynomial vector for these permutations. The designers further claim that polynomial models with many intermediate variables do not improve upon the few variables but high degree polynomial models.
On the other hand, let $I,J\subset P$ be zero-dimensional ideals, then elementary we have that
$I\subset J\qquad\Rightarrow\qquad\mathcal{Z}\left(J\right)\subset\mathcal{Z}\left(I\right),$ (68)
and by the isomorphism theorem for rings
$I\subset J\qquad\Rightarrow\qquad\big{(}P/I\big{)}\big{/}(J/I)\cong P/J.$ (69)
In particular, the latter isomorphism also induces a surjection of finite dimensional $K$-vector spaces $P/I\twoheadrightarrow P/J$ which implies that $\dim_{K}\left(I\right)\geq\dim_{K}\left(J\right)$. In cryptographic terms this means that if $>$-Gröbner bases for $I$ and $J$ are known, then solving for the zero locus of $I$ is at least as hard as solving for the zero locus of $J$ if one uses only generic algorithms.
Back to Ciminion, the designers proposed to study the model in two variables for the keys for the full function, we call this model the fully substituted model. Trivially, we have that
(fully substituted Ciminion model)(iterated Ciminion $(\text{fully substituted {Ciminion} model})\subset(\text{iterated {Ciminion} model}).$ (70)
Since we have a DRL Gröbner basis for the iterated model at hand, we reject the claim by the Ciminion designers that introducing intermediate state variables does not improve upon the fully substituted model. So, the only non-trivial step in a fully determined Ciminion polynomial system is the construction of a univariate polynomial.
As outlined in Corollary 3.2 and Section 3.2, the univariate LEX polynomial can be constructed by iterated substitution of a univariate polynomial into a quadratic polynomial. By [HvdH22] the complexity of univariate polynomial multiplication over a finite field is
$\mathcal{O}\Big{(}d\cdot\log(q)\cdot\log\big{(}d\cdot\log(q)\big{)}\Big{)},$ (71)
where $d$ is the maximum of the degrees of the polynomials. Then, the complexity of constructing the univariate LEX polynomial of Ciminion is bounded by
$\mathcal{O}\Big{(}2^{r_{C}+r_{E}-2}\cdot\log(q)\cdot\log\big{(}2^{r_{C}+r_{E}-2}\cdot\log(q)\big{)}\Big{)}.$ (72)
In Table 3 we present the complexity of term order conversion via fast polynomial multiplication and the GCD with the field equation (Equation 1) for Ciminion. Since the complexities of polynomial multiplication and the GCD are quasi-linear in the degree, one requires more than $110$ rounds for standard Ciminion to achieve $128$ bits of security.
Recall that we proposed a slightly modified variant Ciminion2 (Section 3.3.1) with a key addition to the non-linear component in the last rounds of $p_{C}$ every round of $p_{E}$. As discussed in Section 3.2, this modification breaks the construction of the univariate LEX polynomial via fast polynomial multiplication. For Ciminion2 we therefore either have to do term order conversion to LEX or use the Eigenvalue method. For term order conversion via the probabilistic FGLM algorithm [FGHR14], since we cannot prove $x$-shape position we have to compute the LEX Gröbner basis for the radical ideal which has complexity
$\tilde{\mathcal{O}}\left((r_{C}+r_{E}-1)\cdot 2^{\omega\cdot(r_{C}+r_{E}-1)}\right).$
On the other hand, for Ciminion2 it suffices to solve for the key variables via the Eigenvalue method. Since the Ciminion2 DRL Gröbner basis satisfies the assumption of Section 5, the complexity of characteristic polynomial construction is
$\mathcal{O}\left(2\cdot 2^{\omega\cdot(r_{C}+r_{E}-2)}\right).$ (73)
In Table 3 we present the complexity of the Eigenvalue Method and the GCD with the field equation for Ciminion2. Since the Eigenvalue Method adds a factor $\omega$ (in logarithmic terms), more than $66$ rounds for Ciminion2 already achieve $128$ bits of security.
If we add additional polynomials to the Ciminion or Ciminion2 polynomial systems, then we can recompute the DRL Gröbner bases. It is clear from the Macaulay bound (Theorem 2.5) that a polynomial $f\in P$ with $\deg\left(f\right)\geq 2$ can only increase the Macaulay bound. On the other hand, if $\deg\left(f\right)=1$, then we can eliminate one variable for free. Note that this does not affect the property of being in generic coordinates, but we would have to recompute the Ciminion Macaulay bound (Corollary 3.2). For the former case, we use
$\mathcal{O}\left(\binom{2\cdot(r_{C}+r_{E}-1)}{r_{C}+r_{E}}^{\omega}\right)$ (74)
as baseline for the complexity of DRL Gröbner basis computations. Based on our proven DRL complexity estimate, more than $36$ rounds for Ciminion and Ciminion2 polynomial systems achieve $128$ bits of security.
Lastly, in Table 3 we provide an additional column to compare our estimations to the rational of the Ciminion designers. In [DGGK21, §4.4] it is proposed to model the Ciminion function via two polynomials in two variables, i.e. the fully substituted model, of degree $2^{r_{C}+r_{E}-1}$. Under the assumption that these polynomials are regular we can then use the binomial term in Equation 67 as baseline for DRL Gröbner basis computations
$\mathcal{O}\left(\binom{2+2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1-1}{2\cdot\left(2^{r_{C}+r_{E}-1}-1\right)+1}^{\omega}\right)=\mathcal{O}\left(2^{\omega\cdot(r_{C}+r_{E})}\right),$ (75)
For completeness, this complexity should not be compared to Equation 74. Since our DRL Gröbner basis complexity assumes that we have a Ciminion DRL Gröbner basis at hand plus some additional information that can be modeled via a non-linear polynomial. The complexity of Equation 75 quantifies the complexity of finding a DRL Gröbner basis in the first place without any additional information. Now ignorance of the iterated polynomial model in [DGGK21] becomes even more damaging. After all, we have constructed an iterated DRL Gröbner basis in $\mathcal{O}\left(r_{C}+r_{E}\right)$ while for the fully substituted model we obtain an exponential complexity estimate for finding the DRL basis. Moreover, the strategies for direct polynomial system solving outperform the complexity of finding the DRL basis for the fully substituted model.

### v4

@@ 4.3 Implementation
$\operatorname{rank}\left(\tilde{\mathcal{L}}_{1,8}^{(3)},\dots,\tilde{\mathcal{L}}_{1,8}^{(r_{\mathcal{H}})},\tilde{\mathcal{L}}_{2,8}^{(1)},\dots,\tilde{\mathcal{L}}_{2,8}^{(r_{\mathcal{H}})}\right)=2\cdot r_{\mathcal{H}}-2.$ (56)
So, before the elimination of $62\cdot 7\cdot r_{\mathcal{H}}+$ 6 variables our change of coordinates corresponds to, see Equations 48, 49 and 50,
@@
$\hat{x}_{i}$ $=\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{1,l}^{(i-1)},\qquad$ $3\leq i\leq r_{\mathcal{H}},$ (57)
@@
$\hat{x}_{i+r_{\mathcal{H}}-2}$ $=\sum_{l=1}^{8}(-1)^{\left\lfloor\frac{l-1}{4}\right\rfloor}\cdot x_{2,l}^{(j-1)},\qquad$ $1\leq i\leq r_{\mathcal{H}}.$ (58)
We also note that we did not find a round number $r_{\mathcal{H}}\geq 2$ for which the Gröbner basis extraction fails.
4.4 Cryptanalysis
Next we discuss the cryptanalytic impact of our theoretical analysis. Throughout, this section we will always assume that we can perform the linear change of coordinates from Section 4.2. In particular, we stress that for the Hydra instance from Example 4.4 we are always able to construct a DRL Gröbner basis. Therefore, we have a quadratic Hydra DRL Gröbner basis in $22\cdot r_{\mathcal{H}}-$ 2 equations and variables together with four additional quadratic polynomials at hand. This opens the door for two different attack strategies:
  1. (A)
    Recomputation of the DRL Gröbner basis for the overdetermined system of $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials.
  2. (B)
    Polynomial system solving techniques for the DRL Gröbner basis of $22\cdot r_{\mathcal{H}}-$ 2 quadratic polynomials.
For comparison, we also review the complexity estimation of the designers.
4.4.1 Estimation of the Designers
The Hydra designers considered the iterated polynomial model, see [GØSW23, §7.4] and [GØSW22, Appendix I.2], and reduced it to $22\cdot r_{\mathcal{H}}+$ 2 quadratic polynomials in $22\cdot r_{\mathcal{H}}-$ 2 variables. The designers assumed that this downsized system is semi-regular [BFS04]. Then, the degree of regularity $d_{\operatorname{reg}}$ is given by the index of the first negative coefficient in the power series
$H(t)=\frac{(1-t^{2})^{2\cdot r_{\mathcal{H}}+2}}{(1-t)^{2\cdot r_{\mathcal{H}}-2}},$ (59)
and the complexity of computing the DRL Gröbner basis is then bounded by [BFS04, Theorem 7]
$\mathcal{O}\left(\binom{2\cdot r_{\mathcal{H}}-2+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right).$ (60)
For $r_{\mathcal{H}}^{\ast}=29$ the designers estimate $d_{\operatorname{reg}}=23$, with $\omega=2$ this yields a complexity estimate of $\approx 2^{130.8}$ bits. Moreover, it is claimed that $r_{\mathcal{H}}=r_{\mathcal{H}}^{\ast}+2=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks.
4.4.2 Recomputing the DRL Gröbner Basis
After constructing the Hydra DRL Gröbner basis we have one big advantage compared to the designer estimation: We yield a $\mathbb{F}_{q}$-Boolean polynomial system, see Section 2.4. Thus, in the Macaulay matrix we only have to consider square-free monomials. Effectively, this saves time and space.
Proven Solving Degree Bound.
Recall from Corollary 4.2 that $\operatorname{sd}_{DRL}\left(\mathcal{G}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}$. So, in the worst case we have to construct the full $\mathbb{F}_{q}$-Boolean Macaulay matrix, see Theorem 2.10. The complexity for constructing the $\mathbb{F}_{q}$-Boolean Macaulay matrix was given in Corollary 2.12, for the Hydra parameters $\left|F\right|=4$, $n,d=2\cdot r_{\mathcal{H}}-2$ and $D=2$ we yield the complexity
$\mathcal{O}\left(4\cdot r_{\mathcal{H}}\cdot(2\cdot r_{\mathcal{H}}-1)\cdot\sum_{i=0}^{2\cdot r_{\mathcal{H}}-2}\binom{2\cdot r_{\mathcal{H}}+i}{i+2}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right).$ (61)
Further, by Theorem 2.10 Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrix can be performed in $\mathcal{O}\left(4\cdot 2^{\omega\cdot(2\cdot r_{\mathcal{H}}-2)}\right)$ field operations.
In Table 4 we evaluated these complexities for sample round numbers.
Semi-Regularity Assumption.
If we assume that the Hydra polynomial system is semi-regular after the change of variables (Section 4.2), then Equation 61 improves to
$\mathcal{O}\left(4\cdot r_{\mathcal{H}}\cdot(2\cdot r_{\mathcal{H}}-1)\cdot\sum_{i=0}^{d_{\operatorname{reg}}-2}\binom{2\cdot r_{\mathcal{H}}-2}{i}\cdot\binom{2\cdot r_{\mathcal{H}}+i}{i+2}\right),$ (62)
and the complexity for Gaussian elimination improves to
$\mathcal{O}\left(4\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}-2}\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)^{\omega-1}\right).$ (63)
In Table 4 we evaluated the complexities together with the one from the designers (Section 4.4.1). Construction of the $\mathbb{F}_{q}$-Boolean Macaulay matrix dominates this approach. As expected, under the semi-regularity assumption the Boolean approach is slightly more performative than the designers’ estimation. In particular, an ideal adversary with $\omega=2$ could break up to $r_{\mathcal{H}}=29$ rounds below $128$ bits. Moreover, if the $\mathbb{F}_{q}$-Boolean Macaulay matrix can be constructed with negligible cost, then an ideal adversary could break up to $r_{\mathcal{H}}=34$ rounds.
Table 4: Hydra complexity estimations for recomputing the DRL Gröbner basis. The column $d_{\operatorname{reg}}$ shows the degree of regularity under the semi-regularity assumption. All estimations use $\omega=2$.